### v1

@@ 2.4 Toric elliptic fibrations
We are interested here in understanding elliptic fibrations that are explicit in the structure of a toric hypersurface Calabi-Yau threefold through torically fibered polytopes Avram:1996pj; Kreuzer:1997zg; Skarke:1998. A 4D reflexive polytope $\nabla$ is fibered by a 2D reflexive subpolytope $\nabla_{2}$ when $\nabla_{2}$ is contained in $\nabla$ along a 2D sublattice containing the origin. It was argued in Huang-WT-fibers that any polytope fibration of this kind can be associated with at least one triangulation giving a toric cone structure that is compatible with the toric morphism that projects out the 2D reflexive subpolytope, and hence provides an elliptic fibration structure for the resulting Calabi-Yau threefold. Note that when the toric base onto which this fibration is projected contains curves of self-intersection -3 or below, this triangulation is in general not a standard (i.e. fine, regular, and star) triangulation, but rather in the class studied in Berglund:2016nvh, i.e. of a type that is often referred to as a “vex” triangulation. The standard analysis of Batyrev Batyrev:1993oya generalizes straightforwardly to this kind of triangulation. Note, however, that in most of the analysis of this paper, we do not deal with explicit triangulations; rather, we reformulate the hypersurface defining equation (4) in terms of a Weierstrass model, relying implicitly on the existence of a triangulation compatible with the toric morphism.
The relationship between toric structure and Weierstrass model is easiest to illustrate in the case of a “standard stacking” with the fiber $\mathbb{P}^{2,3,1}$ (fiber $F_{10}$), associated with the generic elliptic fibration over a given base $B_{2}$, where $B_{2}$ is described as a 2D toric variety having rays $v_{i}$. The details of this correspondence and further references are spelled out in detail in Candelas:1997eh; Skarke:1998; Huang-WT-2018-Long; we summarize the story very briefly here. In the simplest case, where the base is weak Fano (and therefore also one of the 16 polytopes in Figure 1), the polytope is defined by the vertices (1, 0; 0, 0), (0, 1; 0, 0) and $\{(-2,-3;v_{i}),v_{i}\in B_{2}\}$. In this situation, the defining equation (4) becomes the long Weierstrass form, sometimes also called Tate form,
@@
where $x,y,z$ are coordinates on $\mathbb{P}^{2,3,1}$ and the $a_{i}$ are determined by the sets of monomials $(j,k;m)$ in the dual polytope $\nabla^{\circ}$; for instance, points $(0,0;m)$ correspond to monomials in $a_{1}$, points (1,1;m)(1,-1;m) correspond to monomials in $a_{2}$, and so on. Note that in this correspondence, a key insight is that the existence of a subpolytope $\nabla_{2}\subset\nabla$ corresponds in the dual polytope $\nabla^{\circ}$ to a projection to the 2D polytope that is dual to the fiber, $\nabla^{\circ}\rightarrow\nabla^{\circ}_{2}$. Building on this basic construction, related polytopes associated with tuned gauge group factors can be realized by removing monomials from the $a_{i}$ so that these line bundle sections vanish to various degrees over the base toric divisors (so-called “Tate tunings,” tables of which with increasing precision can be found in BershadskyEtAlSingularities; Grassi:2011hq; Huang-WT-2018-Long; this corresponds in the original polytope $\nabla$ to adding points associated with a “top” characterizing the associated gauge group. Similarly, for non-weak Fano bases, there is an analogous procedure but to get a reflexive polytope one must take the “dual of the dual” of this construction, [^2] To be precise, stacking $F_{10}$ over a non-Fano base yields a non-reflexive polytope, whose dual contains rational (rather than integral) points; taking the dual of only the integral points yields a reflexive enhancement of the initial polytope, which is the correct analogue of the standard stacking for non-weak-Fano bases. which automatically includes additional points in $\nabla$ associated with non-Higgsable (rigid) gauge factors Morrison:2012np lying over curves in the base of self-intersection -3 or below.
@@
Nonabelian gauge factors in standard stackings of $F_{10}$ are well understood through Tate tunings or equivalently through the study of polytope tops. However, for other fibers besides $F_{10}$, or indeed for nonstandard stackings of $F_{10}$, these analyses fall short. To use tops, for instance, one would need to derive new tops for each fiber type. Instead, we find it convenient to use a more unified framework, in which the different fiber types can all be analyzed with just the Kodaira classification in Table 1. We turn to this more detailed analysis in §2.6 after first briefly discussing genus-one fibrations.
@@ 2.5 Genus-one, elliptic, and Jacobian fibrations
Genus-one fibrations with $k$-sections are related in F-theory to discrete abelian gauge groups. In general, genus-one fibrations with $k$-sections live in Tate-Shafarevich or Weil-Chatalet groups $\mathbb{Z}_{k}$ that share a common $\tau$ function describing the fibration locally, but with different monodromy structure; the Jacobian fibration, corresponding to the identity element of the group, is a true elliptic fibration. F-theory on the Jacobian fibration gives a 6D supergravity theory with gauge group $\mathbb{Z}_{k}$; compactification of this theory on a circle $S^{1}$ with a nontrivial Wilson line gives a 5D theory that can be described as M-theory compactified on the other elements of the Tate-Shafarevich group Braun:2014oya; Morrison:2014era. To understand general elliptic and genus-one fibered Calabi-Yau threefolds with any fiber $F_{n}$ in a parallel fashion to the description above of $F_{10}$ standard stackings with tops and Tate models, we follow a standard approach to the construction of Weierstrass models for the Jacobian fibration of any genus-one or elliptically fibered CY3.
@@ 2.6 Weierstrass models and Nagell’s algorithm
Any elliptic curve $E$ over a field $K$ can be written in long Weierstrass form (5). If the characteristic of $K$ is not 2 or 3444We work in characteristic zero throughout. For recent work on F-theory and related string vacua in positive characteristic, see e.g. Candelas:2019llw; Kachru:2020sio; Kachru:2020abh; Candelas:2023yrg., we can complete the square in $x$ and complete the cube in $y$ to pass to (short) Weierstrass form,
@@
There exists an algorithm, called Nagell’s algorithm, for rewriting an arbitrary elliptic curve $E$, equipped with a zero point $p_{0}$, in Weierstrass form; this algorithm is implemented in e.g. Sage sagemath.
The same procedure of going from an arbitrary defining polynomial to short Weierstrass form can be carried out for an elliptic fibration (with section); this is because elliptic fibrations can equivalently be viewed as elliptic curves whose field of definition is the function field of the base $B_{2}$. The basic idea, as described succinctly in Braun:2014oya, following Deligne, is that given a rational section $z$ of ${\cal O}(-K_{B_{2}})$, one can use the Riemann-Roch theorem to compute the numbers of independent sections of multiples of $K_{B_{2}}$; the number of polynomials in the generating sections at degrees below 6 exceed the number of sections at this degree, and a linear relation follows, which is the Weierstrass equation. Nagell’s algorithm depends crucially on the choice of $p_{0}$, and hence does not apply if $E$ does not have a point over $K$, i.e. if $E$ is a genus-zero curve rather than an elliptic curve; however, in the genus-one context, it is quite natural to instead apply Nagell’s algorithm to the Jacobian $J(E)$. This is done explicitly in an2001jacobians for the cases of 2-, 3-, and 4-sections. To analyze fibrations of the 16 toric fibers in a uniform way, therefore, one can use Nagell’s algorithm to write the defining polynomial of the fibration in Weierstrass form, where the zeroes of the discriminant can be read off straightforwardly from Eq. 8a. This strategy was used in Klevers:2014bqa to give explicit Weierstrass forms for elliptic and (the Jacobians of) genus-one fibrations for all 16 of the 2D toric fiber types.
@@
where $\tilde{\Delta}$ is a generic polynomial in $\alpha,\beta,a_{i}$. Importantly, neither $\tilde{\Delta}$ nor the Weierstrass coefficients $f,g$ vanish automatically when $\alpha$ or $\beta$ do. Thus, on $\alpha=0$ we have $\mathrm{ord}{(f,g,\Delta)}=(0,0,3)$; comparing to Table 1, we see that this corresponds to an $I_{3}$ singularity, and accordingly we expect an $SU(3)$ gauge group whenever $\alpha$ can vanish, i.e. when it consists of more than one monomial. Similarly, on the $\beta=0$ locus we have $\mathrm{ord}{(f,g,\Delta)}=(0,0,2)$, and we expect an $SU(2)$ gauge group. This is the origin of the “generic” $SU(2)\times SU(3)$ gauge group listed for $F_{10}$ in Table 2, as originally identified in Klevers:2014bqa. Note that, as in the discussion of §2.4, when the fiber $F_{10}$ is put in standard form in the first two coordinates of the polytope, the monomials in $\alpha,\beta$ correspond to points in the dual lattice of the form (1,1;m)(-1,1;m) and (2,1;m)(2,-1;m), respectively.
A similar analysis applies to the remaining fiber types. In general, the set of monomials over the dual 2D polytope to a given fiber type will automatically give rise to some additional structure that can include nonabelian gauge factors, U(1) factors, and discrete factors. A full list of the properties of the 16 fiber types, including whether there is a section and the generic gauge group, is given in Table 2. For more details on the specific Weierstrass models for each fiber, see Appendix A.
Polytope Dual $n_{v}$ $n_{p}$ Section? Weierstrass model Gauge group Klevers:2014bqa
@@
$F_{5}$ $F_{12}$ 5 6 Yes 139, $a_{030}=a_{003}=0$ $U(1)^{2}$
$F_{6}$ $F_{11}$ 4 6 Yes 139, $a_{012}=a_{003}=0$ $U(1)\times SU(2)$
$F_{7}$ $F_{7}$ 6 7 Yes 139, $a_{300}=a_{030}=a_{003}=0$ $U(1)^{3}$
$F_{8}$ $F_{8}$ 4 7 Yes 139, $a_{030}=a_{012}=a_{003}=0$ $SU(2)^{2}\times U(1)$
$F_{9}$ $F_{9}$ 5 7 Yes 139, $a_{030}=a_{102}=a_{003}=0$ $SU(2)\times U(1)^{2}$
$F_{10}$ $F_{10}$ 3 7 Yes 145 $SU(2)\times SU(3)$
$F_{11}$ $F_{6}$ 4 8 Yes 145, $a_{6}$ = 0 $(SU(2)\times SU(3)\times U(1))/\mathbb{Z}_{6}$
$F_{12}$ $F_{5}$ 5 8 Yes 139, $a_{120}=a_{030}=a_{102}=$ $SU(2)^{2}\times U(1)^{2}$
@@
Table 2: Properties of the 16 two-dimensional reflexive polytopes and “generic” elliptic fibrations with each fiber. For each polytope, we give its dual polytope, the numbers $n_{v}$ and $n_{p}$ of vertices and points, respectively, whether the generic elliptic fibration admits a section or only a multi-section, the Weierstrass model of the elliptic fibration, and the generic gauge group, as computed in Klevers:2014bqa. Note that some of these gauge groups may have a quotient by a discrete center, not indicated here.
@@
We have described so far various aspects of the geometry of elliptically fibered CY3s. The physics of F-theory essentially provides a dictionary that allows us to relate geometric features of elliptic CY3s to aspects of the 6D supergravity theories that one obtains by compactifying F-theory on these geometric spaces. In this subsection, we will describe this relationship in a little more detail; for a much more detailed review, see e.g. Weigand:2018rez. The material in this subsection is particularly relevant for some of the specific examples that we explore in §5.
@@
Physically, such a compactification gives rise to a 6D supergravity theory with $(1,0)$ supersymmetry. The basic building blocks for such a theory are supergravity multiplets, which include the gravity multiplet, vector multiplets, tensor multiplets, and hypermultiplets. There is always one gravity multiplet, and we write the numbers of the other multiplets in the effective theory as $V$, $T$, and $H$, respectively. As usual in string compactifications, these numbers are determined by the topology of the compactification manifold. For instance, the number of tensor multiplets is fixed straightforwardly by the Hodge numbers of the base:
@@
The gauge group dimension is related to the topological data of the fibration through the Shioda-Tate-Wazir (STW) formula wazir2004arithmetic,
$h^{1,1}(X)=h^{1,1}(B)+1+\operatorname{rk}\operatorname{MW(X)}+\operatorname{rk}G_{\text{nonabelian}},$ (13)
which, geometrically, states that a basis of divisors for the threefold can be formed from the divisors on the base, the zero-section, Mordell-Weil sections, and the components of fibral divisors giving rise to Cartan elements of a nonabelian gauge group.
We next turn to the matter content of the theory, which is controlled by codimension two singularities in the elliptic fibration, as mentioned above (and in some cases by the topology of the divisor supporting the nonabelian gauge group). We can classify the hypermultiplets of the theory by whether they are charged or neutral under the gauge group. The total number $H$ of hypermultiplets and the numbers $H_{\text{neutral}}$ and $H_{\text{charged}}$ of neutral and charged hypermultiplets, respectively, are controlled by the topological data of the fibration and the gauge group by a powerful set of gravitational, gauge-gravitational, and pure gauge anomaly cancellation constraints GreenSchwarzWest6DAnom. The simplest constraints are those that involve only the numbers of charged and uncharged matter fields:
@@
The anomaly constraints on charged matter fields are slightly more intricate, and play a key role in the F-theory/geometry dictionary. For a simple nonabelian gauge group $G$, these can be written as Kumar:2010ru
@@
For a given geometry, these anomaly equations are almost completely sufficient to determine the matter content of a given theory. For a given gauge factor $G$, knowing the anomaly coefficients $a,b$ fixes a unique set of generic BershadskyEtAlSingularities; Taylor:2019ots matter representations for hypermultiplet matter fields in the theory. For example, for SU($N$), generic matter fields are those in the fundamental, two-index antisymmetric, and adjoint representations. Similarly, generic matter charged under a pair of gauge factors $G_{i}\times G_{j}$ arises at the intersection point between the divisors $b_{i},b_{j}$, such as SU($N$) $\times$ SU($M$) bifundamental fields $(\mathbf{N},\bar{\mathbf{M}})$. While more exotic representations can arise in special contexts (see, e.g., Klevers:2017aku), in this paper we encounter only constructions with generic matter content. Tables of the generic matter fields associated with specific gauge group factors (and pairs of factors) on base curves with given intersection numbers can be found in, e.g., Johnson:2016qar.
For abelian factors, a similar set of anomaly equations to those above constrains the matter content (see, e.g., Erler6DAnom; Park:2011wv; ParkAnomalies); in this case, for a single U(1) factor, generic matter has charges $\pm 1$ and $\pm 2$. Abelian factors have an anomaly coefficient $b$ that plays an analogous role to that of nonabelian factors; for multiple abelian factors, this becomes a two-index object $b_{ij}$. In the cases we encounter here, where the abelian factors arise from sections associated with -1 curves in the fiber, the anomaly coefficient for U(1) factors is essentially the canonical class $K$, and we can think of the matter content as simply that of a U(1) factor that arises from an SU(2) factor tuned on the genus-one divisor K-K, broken on the single associated adjoint field.
Finally, there can be a discrete gauge group in the theory. This is harder to see directly from the structure of an elliptic CY3, but ties into the role of genus-one fibered Calabi-Yau threefolds. Physically, as discussed above, genus-one fibered threefolds have a multisection, and live in a group $\mathbb{Z}_{k}$ containing the Jacobian fibration, a genuine elliptic fibration; they are associated with the geometry of F-theory models compactified to 5D with a Wilson line that breaks the discrete gauge group. We will not go deeply into this structure here, but see Braun:2014oya; Morrison:2014era; Anderson:2023wkr; Anderson:2023tfy for more on this subject.

### v2

@@ 2.4 Toric elliptic fibrations
We are interested here in understanding elliptic fibrations that are explicit in the structure of a toric hypersurface Calabi-Yau threefold through torically fibered polytopes Avram et al. (1997); Kreuzer and Skarke (1998); Skarke (1999). A 4D reflexive polytope $\nabla$ is fibered by a 2D reflexive subpolytope $\nabla_{2}$ when $\nabla_{2}$ is contained in $\nabla$ along a 2D sublattice containing the origin. It was argued in Huang and Taylor (2020) that any polytope fibration of this kind can be associated with at least one triangulation giving a toric cone structure that is compatible with the toric morphism that projects out the 2D reflexive subpolytope, and hence provides an elliptic fibration structure for the resulting Calabi-Yau threefold. Note that when the toric base onto which this fibration is projected contains curves of self-intersection -3 or below, this triangulation is in general not a standard (i.e. fine, regular, and star) triangulation, but rather in the class studied in Berglund and Hubsch (2018), i.e. of a type that is often referred to as a “vex” triangulation. The standard analysis of Batyrev Batyrev (1994) generalizes straightforwardly to this kind of triangulation MacFadden and Sheridan (2025). Note, however, that in most of the analysis of this paper, we do not deal with explicit triangulations; rather, we reformulate the hypersurface defining equation (5) in terms of a Weierstrass model, relying implicitly on the existence of a triangulation compatible with the toric morphism.
The relationship between toric structure and Weierstrass model is easiest to illustrate in the case of a “standard stacking” with the fiber $\mathbb{P}^{2,3,1}$ (fiber $F_{10}$), associated with the generic elliptic fibration over a given base $B_{2}$, where $B_{2}$ is described as a 2D toric variety having rays $v_{i}$. The details of this correspondence and further references are spelled out in detail in Candelas et al. (1997); Skarke (1999); Huang and Taylor (2019a); we summarize the story very briefly here. In the simplest case, where the base is weak Fano (and therefore also one of the 16 polytopes in Figure 1), the polytope is defined by the vertices (1, 0; 0, 0), (0, 1; 0, 0) and $\{(-2,-3;v_{i}),v_{i}\in B_{2}\}$. In this situation, the defining equation (5) becomes the long Weierstrass form, sometimes also called Tate form,
@@
where $x,y,z$ are coordinates on $\mathbb{P}^{2,3,1}$ and the $a_{i}$ are determined by the sets of monomials $(j,k;m)$ in the dual polytope $\nabla^{\circ}$; for instance, points $(0,0;m)$ correspond to monomials in $a_{1}$, points (1,1;m)(1,-1;m) correspond to monomials in $a_{2}$, and so on. Note that in this correspondence, a key insight is that the existence of a subpolytope $\nabla_{2}\subset\nabla$ corresponds in the dual polytope $\nabla^{\circ}$ to a projection to the 2D polytope that is dual to the fiber, $\nabla^{\circ}\rightarrow\nabla^{\circ}_{2}$. Building on this basic construction, related polytopes associated with tuned gauge group factors can be realized by removing monomials from the $a_{i}$ so that these line bundle sections vanish to various degrees over the base toric divisors (so-called “Tate tunings”), tables of which with increasing precision can be found in Bershadsky et al. (1996); Grassi and Morrison (2012); Huang and Taylor (2019a); this corresponds in the original polytope $\nabla$ to adding points associated with a “top” characterizing the associated gauge group. Similarly, for non-weak Fano bases, there is an analogous procedure but to get a reflexive polytope one must take the “dual of the dual” of this construction, [^2] To be precise, stacking $F_{10}$ over a non-Fano base yields a non-reflexive polytope, whose dual contains rational (rather than integral) points; taking the dual of only the integral points yields a reflexive enhancement of the initial polytope, which is the correct analogue of the standard stacking for non-weak-Fano bases. which automatically includes additional points in $\nabla$ associated with non-Higgsable (rigid) gauge factors Morrison and Taylor (2012a) lying over curves in the base of self-intersection -3 or below.
@@
Nonabelian gauge factors in standard stackings of $F_{10}$ are well understood through Tate tunings or equivalently through the study of polytope tops. However, for other fibers besides $F_{10}$, or indeed for nonstandard stackings of $F_{10}$, these analyses fall short. To use tops, for instance, one would need to use distinct tops for each fiber type; some progress towards such a geometric analysis of tops for each of the 16 fiber types is given in Bouchard and Skarke (2003). Instead of working explicitly with the geometry of tops, we find it convenient to use a more unified framework, in which the different fiber types can all be analyzed with just the Kodaira classification in Table 1. We turn to this more detailed analysis in §2.6 after first briefly discussing genus-one fibrations.
@@ 2.5 Genus-one, elliptic, and Jacobian fibrations
Genus-one fibrations with $k$-sections are related in F-theory to discrete abelian gauge groups. In general, genus-one fibrations with $k$-sections live in Tate-Shafarevich or Weil-Chatalet groups $\mathbb{Z}_{k}$ that share a common $\tau$ function describing the fibration locally, but with different monodromy structure; the Jacobian fibration, corresponding to the identity element of the group, is a true elliptic fibration. F-theory on the Jacobian fibration gives a 6D supergravity theory with gauge group $\mathbb{Z}_{k}$; compactification of this theory on a circle $S^{1}$ with a nontrivial Wilson line gives a 5D theory that can be described as M-theory compactified on the other elements of the Tate-Shafarevich group Braun and Morrison (2014); Morrison and Taylor (2014). To understand general elliptic and genus-one fibered Calabi-Yau threefolds with any fiber $F_{n}$ in a parallel fashion to the description above of $F_{10}$ standard stackings with tops and Tate models, we follow a standard approach to the construction of Weierstrass models for the Jacobian fibration of any genus-one or elliptically fibered CY3.
@@ 2.6 Weierstrass models and Nagell’s algorithm
Any elliptic curve $E$ over a field $K$ can be written in long Weierstrass form (6). If the characteristic of $K$ is not 2 or 3444We work in characteristic zero throughout. For recent work on F-theory and related string vacua in positive characteristic, see e.g. Candelas et al. (2020); Kachru et al. (2024, 2020); Candelas et al. (2023)., we can complete the square in $x$ and complete the cube in $y$ to pass to (short) Weierstrass form,
@@
There exists an algorithm, called Nagell’s algorithm, for rewriting an arbitrary elliptic curve $E$, equipped with a zero point $p_{0}$, in Weierstrass form; this algorithm is implemented in e.g. Sage The Sage Developers (2024).
The same procedure of going from an arbitrary defining polynomial to short Weierstrass form can be carried out for an elliptic fibration (with section); this is because elliptic fibrations can equivalently be viewed as elliptic curves whose field of definition is the function field of the base $B_{2}$. The basic idea, as described succinctly in Braun and Morrison (2014), following Deligne, is that given a rational section $z$ of ${\cal O}(-K_{B_{2}})$, one can use the Riemann-Roch theorem to compute the numbers of independent sections of multiples of $K_{B_{2}}$; the number of polynomials in the generating sections at degrees below 6 exceed the number of sections at this degree, and a linear relation follows, which is the Weierstrass equation. Nagell’s algorithm depends crucially on the choice of $p_{0}$, and hence does not apply if $E$ does not have a point over $K$, i.e. if $E$ is a genus-zero curve rather than an elliptic curve; however, in the genus-one context, it is quite natural to instead apply Nagell’s algorithm to the Jacobian $J(E)$. This is done explicitly in An et al. (2001) for the cases of 2-, 3-, and 4-sections. To analyze fibrations of the 16 toric fibers in a uniform way, therefore, one can use Nagell’s algorithm to write the defining polynomial of the fibration in Weierstrass form, where the zeroes of the discriminant can be read off straightforwardly from Eq. 9a. This strategy was used in Klevers et al. (2015) to give explicit Weierstrass forms for elliptic and (the Jacobians of) genus-one fibrations for all 16 of the 2D toric fiber types.
@@
where $\tilde{\Delta}$ is a generic polynomial in $\alpha,\beta,a_{i}$. Importantly, neither $\tilde{\Delta}$ nor the Weierstrass coefficients $f,g$ vanish automatically when $\alpha$ or $\beta$ do. Thus, on $\alpha=0$ we have $\mathrm{ord}{(f,g,\Delta)}=(0,0,3)$; comparing to Table 1, we see that this corresponds to an $I_{3}$ singularity, and accordingly we expect an $SU(3)$ gauge group whenever $\alpha$ can vanish, i.e. when it consists of more than one monomial. Similarly, on the $\beta=0$ locus we have $\mathrm{ord}{(f,g,\Delta)}=(0,0,2)$, and we expect an $SU(2)$ gauge group. This is the origin of the “generic” $SU(2)\times SU(3)$ gauge group listed for $F_{10}$ in Table 2, as originally identified in Klevers et al. (2015). Note that, as in the discussion of §2.4, when the fiber $F_{10}$ is put in standard form in the first two coordinates of the polytope, the monomials in $\alpha,\beta$ correspond to points in the dual lattice of the form (1,1;m)(-1,1;m) and (2,1;m)(2,-1;m), respectively.
A similar analysis applies to the remaining fiber types. In general, the set of monomials over the dual 2D polytope to a given fiber type will automatically give rise to some additional structure that can include nonabelian gauge factors, U(1) factors, and discrete factors. A full list of the properties of the 16 fiber types, including whether there is a section and the generic gauge group, is given in Table 2.  [^5] Thanks to Thorsten Schimannek for a helpful discussion on the global structure of the group in the case of fiber $F_{4}$. For more details on the specific Weierstrass models for each fiber, see Appendix A.
Polytope Dual $n_{v}$ $n_{p}$ Section? Weierstrass model Gauge group Klevers et al. (2015)
@@
$F_{5}$ $F_{12}$ 5 6 Yes 140, $a_{030}=a_{003}=0$ $U(1)^{2}$
$F_{6}$ $F_{11}$ 4 6 Yes 140, $a_{012}=a_{003}=0$ $U(1)\times SU(2)$
$F_{7}$ $F_{7}$ 6 7 Yes 140, $a_{300}=a_{030}=a_{003}=0$ $U(1)^{3}$
$F_{8}$ $F_{8}$ 4 7 Yes 140, $a_{030}=a_{012}=a_{003}=0$ $SU(2)^{2}\times U(1)$
$F_{9}$ $F_{9}$ 5 7 Yes 140, $a_{030}=a_{102}=a_{003}=0$ $SU(2)\times U(1)^{2}$
$F_{10}$ $F_{10}$ 3 7 Yes 146 $SU(2)\times SU(3)$
$F_{11}$ $F_{6}$ 4 8 Yes 146, $a_{6}$ = 0 $(SU(2)\times SU(3)\times U(1))/\mathbb{Z}_{6}$
$F_{12}$ $F_{5}$ 5 8 Yes 140, $a_{120}=a_{030}=a_{102}=$ $SU(2)^{2}\times U(1)^{2}$
@@
Table 2: Properties of the 16 two-dimensional reflexive polytopes and “generic” elliptic fibrations with each fiber. For each polytope, we give its dual polytope, the numbers $n_{v}$ and $n_{p}$ of vertices and points, respectively, whether the generic elliptic fibration admits a section or only a multi-section, the Weierstrass model of the elliptic fibration, and the generic gauge group, as computed in Klevers et al. (2015). Note that some of these gauge groups may have a quotient by a discrete center not indicated here; we have included some discrete quotient where known, or where they can be inferred from the matter spectrum.
@@
We have described so far various aspects of the geometry of elliptically fibered CY3s. The physics of F-theory essentially provides a dictionary that allows us to relate geometric features of elliptic CY3s to aspects of the 6D supergravity theories that one obtains by compactifying F-theory on these geometric spaces. In this subsection, we will describe this relationship in a little more detail; for a much more detailed review, see e.g. Weigand (2018). The material in this subsection is particularly relevant for some of the specific examples that we explore in §5.
@@
Physically, such a compactification gives rise to a 6D supergravity theory with $(1,0)$ supersymmetry. The basic building blocks for such a theory are supergravity multiplets, which include the gravity multiplet, vector multiplets, tensor multiplets, and hypermultiplets. There is always one gravity multiplet, and we write the numbers of the other multiplets in the effective theory as $V$, $T$, and $H$, respectively. As usual in string compactifications, these numbers are determined by the topology of the compactification manifold. For instance, in the absence of (4, 6) loci giving SCFT’s, the number of tensor multiplets is fixed straightforwardly by the Hodge numbers of the base:
@@
The gauge group dimension is related to the topological data of the fibration through the Shioda-Tate-Wazir (STW) formula Wazir (2004),
$h^{1,1}(X)=h^{1,1}(B)+1+\operatorname{rk}\operatorname{MW(X)}+\operatorname{rk}G_{\text{nonabelian}}(+\operatorname{rk}\operatorname{SCFT}),$ (14)
which, geometrically, states that a basis of divisors for the threefold can be formed from the divisors on the base, the zero-section, Mordell-Weil sections, and the components of fibral divisors giving rise to Cartan elements of a nonabelian gauge group. Note that we have included in the last term also contributions for singular fibrations with (4, 6) loci corresponding to the number of tensors in the associated SCFT, generalizing the usual STW formula for smooth Calabi-Yau manifolds; in the flop phase where these points in the base are blown up these are incorporated into $h^{1,1}(B^{\prime})$ for the blown up base, and otherwise this contribution corresponds to non-flat fibers.
We next turn to the matter content of the theory, which is controlled by codimension two singularities in the elliptic fibration, as mentioned above (and in some cases by the topology of the divisor supporting the nonabelian gauge group). We can classify the hypermultiplets of the theory by whether they are charged or neutral under the gauge group. The total number $H$ of hypermultiplets and the numbers $H_{\text{neutral}}$ and $H_{\text{charged}}$ of neutral and charged hypermultiplets, respectively, are controlled by the topological data of the fibration and the gauge group by a powerful set of gravitational, gauge-gravitational, and pure gauge anomaly cancellation constraints Green et al. (1985). The simplest constraints are those that involve only the numbers of charged and uncharged matter fields:
@@
The anomaly constraints on charged matter fields are slightly more intricate, and play a key role in the F-theory/geometry dictionary. For a simple nonabelian gauge group $G$, these can be written as Kumar et al. (2010)
@@
For a given geometry, these anomaly equations are almost completely sufficient to determine the matter content of a given theory. For a given gauge factor $G$, knowing the anomaly coefficients $a,b$ fixes a unique set of generic Bershadsky et al. (1996); Taylor and Turner (2019) matter representations for hypermultiplet matter fields in the theory. For example, for SU($N$), generic matter fields are those in the fundamental, two-index antisymmetric, and adjoint representations. Similarly, generic matter charged under a pair of gauge factors $G_{i}\times G_{j}$ arises at the intersection point between the divisors $b_{i},b_{j}$, such as SU($N$) $\times$ SU($M$) bifundamental fields $(\mathbf{N},\bar{\mathbf{M}})$. While more exotic representations can arise in special contexts (see, e.g., Klevers et al. (2017)), in this paper we encounter only constructions with generic matter content. Tables of the generic matter fields associated with specific gauge group factors (and pairs of factors) on base curves with given intersection numbers can be found in, e.g., Johnson and Taylor (2016).
For abelian factors, a similar set of anomaly equations to those above constrains the matter content (see, e.g., Erler (1994); Park and Taylor (2012); Park (2012)); in this case, for a single U(1) factor, generic matter has charges $\pm 1$ and $\pm 2$. Abelian factors have an anomaly coefficient $b$ that plays an analogous role to that of nonabelian factors; for multiple abelian factors, this becomes a two-index object $b_{ij}$. In the cases we encounter here, where the abelian factors arise from sections associated with -1 curves in the fiber, the anomaly coefficient for U(1) factors is essentially the canonical class $K$, and we can think of the matter content as simply that of a U(1) factor that arises from an SU(2) factor tuned on the genus-one divisor K-K, broken on the single associated adjoint field.
Finally, there can be a discrete gauge group in the theory. This is harder to see directly from the structure of an elliptic CY3, but ties into the role of genus-one fibered Calabi-Yau threefolds. Physically, as discussed above, genus-one fibered threefolds have a multisection, and live in a group $\mathbb{Z}_{k}$ containing the Jacobian fibration, a genuine elliptic fibration; they are associated with the geometry of F-theory models compactified to 5D with a Wilson line that breaks the discrete gauge group. We will not go deeply into this structure here, but see Braun and Morrison (2014); Morrison and Taylor (2014); Anderson et al. (2024, 2023) for more on this subject.